# Po-Heng Chou1, Da-Chih Lin2, Hung-Yu Wei2, Walid Saad3, and Yu Tsao1

### v2

@@ V-C Overall Trends Across Horizons
Tables I summarize results for $T_{s}=3$ s. Moving from $T_{p}=1$ s to $T_{p}=2$ s generally improves F1 for most models, indicating that a slightly longer horizon exposes more pre-failure cues without overly diluting short-term dynamics. While gains from $T_{p}=2$ s to $T_{p}=3$ s are modest for several baselines, TimesNet benefits from the longer horizon and reaches its peak performance at $T_{s}=3$ s and $T_{p}=3$ s. Hence, $T_{p}=2$ s remains a favorable operating point for low-latency prediction (e.g., with CNN), whereas $T_{p}=3$ s can yield the highest overall reliability with TimesNet under sufficient context. Moreover, all evaluated models demonstrated excellent discriminative performance, with AUC values exceeding 0.95, suggesting highly accurate prediction of events.
@@
(a) CNN [10] with $T_{p}=2$ s
(b) XGBoost [12] with $T_{p}=2$ s
(c) TimesNet [15] with $T_{p}=3$ s
Figure 3: Prediction hits of the top-3 models based on Tables I. Profiles use $T_{s}=3$ s and $I_{s}=0.1$ s; CNN and XGBoost are shown with $T_{p}=2$ s, and TimesNet with $T_{p}=3$ s. Red markers indicate successful predictions, i.e., time slots where an RLF was predicted to occur within the next $T_{p}$ seconds.
V-D Model-wise Comparison
V-E Effect of Observation Window $T_{s}$
@@
Extending the observation window from $T_{s}=1$ s to $T_{s}=3$ s roughly doubles or triples the theoretical floating point operations per second (FLOPs) for all models. For example, the TimesNet and CNN models respectively grow by about $2.8\times$ and $2.9\times$ in FLOPs. However, the measured inference latency increases by only around $0.2$ ms on our CPU platform, demonstrating that longer temporal contexts can be leveraged with negligible runtime cost and an overall favorable accuracy-latency tradeoff.
V-G Per-horizon Highlights
$T_{p}=1$ s
Short horizons favor recall-oriented models such as CNN and XGBoost, yet their precision remains limited due to scarce pre-failure cues within 1 s.
$T_{p}=2$ s
CNN achieves its best balance between accuracy and latency, reaching F1 = 0.8208 at $T_{s}=3$ s. TimesNet and XGBoost follow closely, showing that moderate prediction horizons ($\approx 2$ s) effectively capture early degradation patterns.
$T_{p}=3$ s
TimesNet achieves the highest overall F1 = 0.8498, benefiting from its multi-period 2D temporal blocks that aggregate longer-term variations. This indicates that under sufficient context length, TimesNet provides the most reliable long-horizon early-warning capability among all tested models.
V-H Operating Points and Thresholding
All reported scores use the F1-optimal threshold $\tau$ chosen on the validation set. In safety-critical operation, $\tau$ can be shifted at deployment to favor recall (earlier alarms) at the expense of precision. We found that model ranking remains stable under reasonable $\tau$ shifts, indicating robust comparative behavior across methods.
V-I Time-domain Early-warning Behavior
Figure 3 visualizes time-indexed prediction hits for the top three models at $T_{s}=3$ s. Red markers cluster before RLF timestamps, confirming that alarms are typically raised within the allowed horizon. Intervals without RLFs remain mostly unflagged, reducing alarm fatigue. Table II further quantifies hit policies relevant to operations: requiring any two points (i.e., two alarms within the horizon) preserves near-perfect coverage for CNN and TimesNet while suppressing sporadic single-point spikes; stricter policies (three consecutive points) reduce false positives but may miss faster-onset events. These trade-offs offer a tunable path from research metrics to actionable policies (e.g., trigger redundancy or selective band locking upon two consecutive positives).
V-J Error Analysis
Qualitative inspection of false negatives reveals two common patterns: (i) rapid-degradation cases where RSRP/RSRQ drop abruptly with minimal prelude, and (ii) interference-driven perturbations that mimic normal variability until the final seconds. False positives frequently coincide with transient neighbor dominance changes that do not culminate in RLF. Incorporating richer neighbor-cell context or selective protocol cues (e.g., imminent reconfiguration signals) can mitigate such errors.
V-K Runtime and Deployability
Short windows ($T_{s}\leq 3$ s at 10 Hz) keep inference budgets low. CNN and TimesNet offer favorable trade-offs between accuracy and latency with regular compute patterns amenable to UEs or edge gateways. In practice, we recommend pairing a recall-biased operating point with a confirmation policy (two consecutive positives) to enable proactive actions while controlling nuisance alarms.
@@ VI Conclusion
This work introduced a measurement-driven framework for early RLF prediction in 5G NSA railway environments. Using 10 Hz metro-train measurements, six learning models were benchmarked across multiple temporal configurations. TimesNet achieved the highest F1 with a 3 s observation window ($T_{s}=3$ s) and a 3 s prediction horizon ($T_{p}=3$ s), while CNN delivered comparable performance with lower latency at $T_{p}=2$ s. The results show that deep temporal models can anticipate reliability degradation several seconds in advance using lightweight radio indicators, enabling actionable early warning for proactive control in railway communications. Future research will extend the proposed framework to multi-cell and multi-train scenarios, where inter-train interference and network-side coordination are jointly considered. Moreover, self-adaptive learning at the edge and transfer learning across different routes and operators will be investigated to maintain prediction robustness under dynamic radio environments while reducing data labeling costs for large-scale deployment.

### v4

@@ V-C Overall Trends Across Horizons
Tables I summarize results for $T_{s}=3$ s. Moving from $T_{p}=1$ s to $T_{p}=2$ s generally improves F1 for most models, indicating that a slightly longer horizon exposes more pre-failure cues without overly diluting short-term dynamics. While gains from $T_{p}=2$ s to $T_{p}=3$ s are modest for several baselines, TimesNet benefits from the longer horizon and reaches its peak performance at $T_{s}=3$ s and $T_{p}=3$ s. Hence, $T_{p}=2$ s remains a favorable operating point for low-latency prediction (e.g., with CNN), whereas $T_{p}=3$ s can yield the highest overall reliability with TimesNet under sufficient context. Moreover, all evaluated models achieved AUC values exceeding 0.95, indicating strong ranking capability, although such ROC-oriented metrics should be interpreted together with PR-AUC and alarm-centric measures under severe class imbalance.
V-D Model-wise Comparison
V-E Effect of Observation Window $T_{s}$
@@
Extending the observation window from $T_{s}=1$ s to $T_{s}=3$ s roughly doubles or triples the theoretical FLOPs for all models. However, the measured inference latency increases by only about $0.2$ ms on our CPU platform, indicating that longer temporal contexts can be exploited with negligible runtime penalty.
V-G Time-domain Early-warning Behavior
Figure 3 visualizes time-indexed prediction hits for the top three models at $T_{s}=3$ s. Red markers cluster before labeled reliability-breakdown timestamps, confirming that alarms are typically raised within the allowed horizon. Intervals without labeled breakdown events remain mostly unflagged, suggesting that nuisance alarms can be partially controlled under suitable trigger policies. Table II further quantifies hit policies relevant to operations: requiring any two points (i.e., two alarms within the horizon) preserves near-perfect coverage for CNN and TimesNet while suppressing sporadic single-point spikes; stricter policies (three consecutive points) reduce false positives but may miss faster-onset events. These trade-offs provide a tunable path from research metrics to deployment-oriented alarm policies, for example, issuing a proactive warning only after two consecutive positive decisions.
@@ V-G Time-domain Early-warning Behavior
V-H Error Analysis
Qualitative inspection shows that false negatives mainly arise from abrupt degradation with little prelude or interference-driven perturbations that resemble normal variability, while false positives often coincide with transient neighbor-dominance changes. These observations motivate confirmation-based trigger policies and false-alarm-budget evaluation in practical railway settings.
@@ VI Conclusion
This work presented a measurement-driven benchmark for early prediction of RLF-related reliability breakdown events in 5G NSA railway environments. Using 10 Hz metro-train measurements, six learning models were evaluated across multiple temporal configurations. TimesNet achieved the best overall F1 at $(T_{s},T_{p})=(3\,\mathrm{s},3\,\mathrm{s})$, while CNN offered a strong latency-reliability trade-off at $T_{p}=2$ s. The results suggest that lightweight device-side radio indicators can support seconds-ahead reliability warning for IoT-enabled transportation systems. Future work will extend the benchmark to multi-line and cross-environment settings and further investigate event-level and false-alarm-budget evaluation.